OpenStudy (anonymous):

What is the solution for x in the proportion two over six equals x over twenty-one? 7 6 4 3

OpenStudy (anonymous):

@Sunshine447

jhonyy9 (jhonyy9):

2/6 = x/21 cross multiplie 42=6x divide both sides by 6 7=x